OpenStudy (anonymous):

Find the type and number of solutions for the equation. x^2 - 4x = -4

OpenStudy (anonymous):

HELP ME PLEASE

OpenStudy (anonymous):

two distinct nonreal complex solutions no solutions two distinct real solutions one distinct real solution

OpenStudy (anonymous):

add 4 in both sides and you can use the formula (A-B)^2=A^2-2AB+B^2

OpenStudy (anonymous):

which is the right one? two distinct nonreal complex solutions no solutions two distinct real solutions one distinct real solution

OpenStudy (anonymous):

(x-2)^2=0 what o you think. How man solutions.

OpenStudy (anonymous):

1?

OpenStudy (anonymous):

that is right:)

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

u r wlcm:)

OpenStudy (shamim):

\[x ^{2}-4*x+4=0 \] or, \[(x-2)^{2}=0\] or\[(x-2)=0\] or,\[x=2\]

OpenStudy (shamim):

so ur equation has only 1 solution

OpenStudy (anonymous):

thanks ^^